OpenStudy (anonymous):

HELP!?! Given its parent function, f(x) = 2x, what is the equation of the function shown?

OpenStudy (cwrw238):

hint: f(x) = 2^x when x = 0 f(x) = 2^0 = 1 so the graph of f(x) = 2^x will pass through +1 on the vertical axis

OpenStudy (anonymous):

Hmmm I think I need further help reallly new to this

OpenStudy (anonymous):

how would I do that/

OpenStudy (anonymous):

on the graph for x=0 then f(x)=2 so substitute 0 in place of x to find 2

OpenStudy (cwrw238):

ok lets try inserting values of x into each option first option: x = o h(x) = 2^0 - 1 = 1-1 = 0 the graph passes thru vertical axis at 2 so not this one

OpenStudy (cwrw238):

second: x = 0, (x) = 1 - 3 = -2 not this one

OpenStudy (cwrw238):

so try 3rd and 4th options - can you do that?

OpenStudy (cwrw238):

if x = 0 fits you'll then need to try x = 1 to check the point (1,5)

OpenStudy (cwrw238):

do you follow me ok?

OpenStudy (anonymous):

I'm trying to thank you

OpenStudy (cwrw238):

correction: * second option: x=o , h(x) = 2^0 -3 = -2

OpenStudy (anonymous):

so what do I do now?

OpenStudy (anonymous):

Ohh you mean it's that one

OpenStudy (cwrw238):

plug in x = 0 into the third option h(x) = 3(2^x) - 1 and see if this gives you 2 (that is to check if the graph passes through h(x) axis at 2)

OpenStudy (cwrw238):

no its not second one because h(x) = -2 not 2

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

it gave me 2

OpenStudy (anonymous):

oh so it's A

OpenStudy (cwrw238):

right - thats ok so far now notice the graph also goes through the point (1,5) that is when x = 1 h(x) = 5 so now plug in x = 1 into 3rd option and check if you get 5

OpenStudy (anonymous):

what option?

OpenStudy (anonymous):

Oh ok I see it now @cwrw238

OpenStudy (anonymous):

i think it's c?